Marsha has $23,479 in the bank. If she deposits another 25% of her total into her account, what percent of the new total must she withdraw to give her the same total she started with? (round the answer to the nearest whole number)
Answer by Theo(13342)


the original amount is 23479.

add 25% to this and you get 23479 + (.25 * 23479) = 29348.75

now you want to know what percent of the new amount do you have to subtract to get back to the original amount.

let x = the rate that you have to multiply the new amount by in order to find out what to subtract.

your new formula is 29348.75 - (x * 29348.75) = 23479

subtract 23479 from both sides of this equation and add x * 29348.75 to both sides of this equation to get:

29348.75 - 23479 = x * 29348.75

combine like terms to get:

5869.75 = x * 29348.75

divide both sides of this equation by 29348.75 to get:

5869.75 / 29348.75 = x

solve for x to get x = .2

.2 is the rate that you have to multiply 29348.75 by to find how much you need to subtract from 29348.75 in order to get back to 23479.

a rate of .2 is the same as a percent of 20.

.2 = 20%.

you would have to subtract 20% of the new amount to get back to the original amount.